A total of tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold. How many adult tickets were sold?
Answer by josgarithmetic(39623)

Assign variables to how many of each ticket.
Account for the number of tickets sold.
Solve for each variable.

Your description means this:
x, the number of student tickets
y, the number of adult tickets
T, the total number of tickets sold
x=2y
x+y=T
2y+y=T
3y=T or y=T/3


That is all that you can do. Information is missing from your description.